OpenStudy (anonymous):

A bottle has a mass of 25 g when empty and 125 g when filled with water. When filled with another liquid, the total mass is 140 g. What is the density of the liquid?

OpenStudy (ivancsc1996):

Well density is the amount of mass there is per volume. In the first case we know that the mass of water added was a hundred grams (we know this from the substraction of the final mass and the original). We can know the volume of the bottle of the bottle by dividing the mass of water by the density of water. In the second case we know that the mass of liquid added was a hundred and fifteen grams. Now to figure out the density of the liquid you just need to divide the mass added by the volume. The answer is 1.15 grams per cubic centimeter.\[\rho=\frac{ m }{ V } \rightarrow V=\frac{ m }{ \rho } \]\[\rho _{water}=1\frac{ g }{ cm ^{3} }\]\[m _{water}=\Delta m=m _{f}-m _{i}=125g-25g=100g\]\[V=\frac{ m }{ V }=\frac{ 100g }{ 1\frac{ g }{ cm ^{3} } }=100cm ^{3}\]\[m _{liquid}=\Delta m=m _{f}-m _{i}=140g-25g=115g\]\[\rho _{liquid}=\frac{ m _{liquid} }{ V }=\frac{ 115g }{ 100cm ^{3}}=1.15\frac{ g }{ cm ^{3} }\]

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

That was an awesome response